# The hypothetical track-length fitting algorithm for energy measurement in liquid argon TPCs

### v2

@@ 6.5.1 Reproduction of data distributions
We attempt to reproduce the data’s distributions by applying scale corrections to the MC simulation sample’s values. The distribution of the MPV ratio between MC simulation and data, as shown in figure 17, is fitted using two linear functions: one flat function for the MIP region and the other with a slope for the Bragg peak region, to derive the scale corrections. It is shown that the two linear functions intersect at a value of 2.39 . Consequently, the scale correction for the Bragg peak region is applied to hits with greater than 2.39 . As a result, figure 18 shows improved agreement in MPV values between data and MC simulation in the Bragg peak region compared to the state before applying the correction, as shown in figure 16. Note that widths of the Gaussian contribution also have better agreements in the Bragg peak region after applying the scale correction, but with non-negligible differences with respect to their statistical uncertainties. For the MIP region, we test two constant scale corrections with 0.985 and 0.975 to consider the discrepancy for from 1.8 to 2.3 shown in figure 17.
@@ 6.5.2 Modified box model parameter uncertainties
Another important point to note here is that shifting the modified box model parameters by one standard deviation can change distributions significantly in the Bragg peak region as shown in figure 19. Therefore, a corresponding systematic uncertainty should be considered for the hypothetical track-length fitting algorithm’s performance results in figure 15 for data. The study shown in Section 6.4 is repeated for data with the modified box model parameters shifted up and down by one standard deviations which gives 8 sets of and besides the central set as shown in Table 2.
@@ 6.5.4 Results
Figures 20 and 21 summarize the hypothetical track length-fitting method’s performance results derived from the studies described above. Energy measurement resolution results are stable. For fractional biases, there are two points to note here. The first point is about scale corrections for the MC sample. While the correction factor for values exceeding 2.39 MeV/cm can reach up to 5%, and the correction for values in the range of 1.8 to 2.3 MeV/cm can reach up to 2.5%, the latter correction has a more significant impact on the fractional bias result. This can be understood from the slope of the Bethe-Bloch formula as a function of residual range shown in figure 1. A few-percent shift in could lead to a larger change in the residual range for small values compared to large values since the slope of the energy loss function is falling down as a function of residual range. It can also explain the tendency shown in central data that charged pions with higher show larger fractional biases in their measured . The second point is that shifts in modified box model parameters for data have significant impact on fractional bias results. These two points conclude that better understanding of measurements leads to smaller systematic uncertainty on the scale of measured by the hypothetical track-length fitting method.
@@ 7 Summary
We introduce the hypothetical track-length fitting method, a particle energy measurement technique which can be used in a liquid argon time projection chamber. Its performance for charged pion energy measurement is studied in detail using secondary charged pions from data collected by the ProtoDUNE-SP detector with a charged pion beam and in a corresponding Monte Carlo sample. For charged pion tracks with kinetic energy () less than 400 MeV and with the Bragg peak signature, the energy resolution is better than 6%, and the fractional bias is less than 5% for MC sample using true as reference. For direct comparison of the performance between the data and the MC samples, range-based is used as a reference after a validation that it can describe the true well for charged pions passing a selection based on . After subtracting the proton contribution from distributions of the fractional energy residual, the resolution is measured to be better than 8% (7%), and the fractional bias is smaller than 8% (3%) for the data (MC simulation) sample. Differences between the data and MC simulation samples are considered in order to calculate the energy scale correction and the systematic uncertainty on the scale and resolution. Additional studies find that the resolution results are stable and the fractional bias results are sensitive to the scale in the MIP region and also to the recombination model. Since the algorithm fits for the expected total track length until the particle would stop for an incomplete track, it can be used to measure the energies of charged pions absorbed by argon nuclei in the detector material. Absorption is the dominant inelastic scattering interaction between charged pions and argon nuclei for charged-pion kinetic energies less than 300 MeV. In addition, the method of using stopping secondary charged pions to characterize the energy measurement performance can be used in any LArTPC by collecting neutrino interaction events with stopping charged pions in the final state.

### v3

@@ 6.5.1 Reproduction of data distributions
We attempt to reproduce the data’s distributions by applying scale corrections to the MC simulation sample’s values. The distribution of the MPV ratio between MC simulation and data, as shown in figure 17(b), is fitted using two linear functions: one flat function for the MIP region and the other with a slope for the Bragg peak region, to derive the scale corrections. It is shown that the two linear functions intersect at a value of 2.39 . Consequently, the scale correction for the Bragg peak region is applied to hits with greater than 2.39 . As a result, figure 18 shows improved agreement in MPV values between data and MC simulation in the Bragg peak region compared to the state before applying the correction, as shown in figure 16. Note that widths of the Gaussian contribution also have better agreements in the Bragg peak region after applying the scale correction, but with non-negligible differences with respect to their statistical uncertainties. For the MIP region, we test two constant scale corrections with 0.985 and 0.975 to consider the discrepancy for from 1.8 to 2.3 shown in figure 17(b).
@@ 6.5.2 Modified box model parameter uncertainties
Another important point to note here is that shifting the modified box model parameters by one standard deviation can change distributions significantly in the Bragg peak region as shown in figure 19. Therefore, a corresponding systematic uncertainty should be considered for the TLEFit algorithm’s performance results in figure 15 for data. The study shown in Section 6.4 is repeated for data with the modified box model parameters shifted up and down by one standard deviations which gives 8 sets of and besides the central set as shown in Table 2. The envelope among the 8 systematic variations is taken as systematic uncertainty for each fractional bias and resolution data point.
@@ 6.5.4 Results
Figures 20 and 21 summarize the track length-fitting method’s performance results derived from the studies described above. Energy measurement resolution results are stable. For fractional biases, there are two points to note here. The first point is about scale corrections for the MC sample. While the correction factor for values exceeding 2.39 MeV/cm can reach up to 5%, and the correction for values in the range of 1.8 to 2.3 MeV/cm can reach up to 2.5%, the latter correction has a more significant impact on the fractional bias result. This can be understood from the slope of the Bethe-Bloch formula as a function of residual range shown in figure 1. A few-percent shift in could lead to a larger change in the residual range for small values compared to large values since the slope of the energy loss function is falling down as a function of residual range. It can also explain the tendency shown in central data that charged pions with higher show larger fractional biases in their measured . The second point is that shifts in modified box model parameters for data have significant impact on fractional bias results. These two points conclude that better understanding of measurements leads to smaller systematic uncertainty on the scale of measured by the TLEFit method.
@@ 7 Summary
We introduce the track-length extension fitting (TLEFit) algorithm for measuring the kinetic energies of inelastically interacting particles in liquid argon time projection chambers. The algorithm’s performance for charged pion energy measurement is studied in detail using secondary charged pions from data collected by the ProtoDUNE-SP detector with a charged pion beam and in a corresponding Monte Carlo sample. For charged pion tracks with kinetic energy () less than 400 MeV and with the Bragg peak signature, the energy resolution is better than 6.5%, and the fractional bias is less than 4% for MC sample using true as reference. For direct comparison of the performance between the data and the MC samples, range-based is used as a reference after a validation that it can describe the true well for charged pions passing a selection based on . After subtracting the proton contribution from distributions of the fractional energy residual, the resolution is measured to be better than 8% (7%), and the fractional bias is smaller than 8% (3%) for the data (MC simulation) sample. Differences between the data and MC simulation samples are considered in order to calculate the energy scale correction and the systematic uncertainty on the scale and resolution. Additional studies find that the resolution results are stable and the fractional bias results are sensitive to the scale in the MIP region and also to the recombination model. Since the algorithm fits for the expected total track length until the particle would stop for an incomplete track, it can be used to measure the energies of charged pions absorbed by argon nuclei in the detector material. Absorption is the dominant inelastic scattering interaction between charged pions and argon nuclei for charged-pion kinetic energies less than 300 MeV. In addition, the method of using stopping secondary charged pions to characterize the energy measurement performance can be used in any LArTPC by collecting neutrino interaction events with stopping charged pions in the final state.